OpenStudy (anonymous):

State the restrictions and simplify: 1/x-2 - 1/3x+4

OpenStudy (anonymous):

\[\frac{ 1 }{ x-2 } - \frac{ 1 }{ 3x-4 }\]

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this

jimthompson5910 (jim_thompson5910):

To subtract the fractions, you need to make sure the denominators are equal.

OpenStudy (anonymous):

I got that and for some reason I am stuck I will draw what I have so you can tell me what I am doing wrong

jimthompson5910 (jim_thompson5910):

alright go for it

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

i think you meant to say 3x-4 instead of 3x+4 right?

jimthompson5910 (jim_thompson5910):

or is it really 3x+4 ?

OpenStudy (anonymous):

no its 3x+4

jimthompson5910 (jim_thompson5910):

ok let me make the fix on my steps then

jimthompson5910 (jim_thompson5910):

what you have is correct Here's what I got \[\Large \frac{ 1 }{ x-2 } - \frac{ 1 }{ 3x+4 }\] \[\Large \frac{ 1(3x+4) }{ (x-2)(3x+4) } - \frac{ 1 }{ 3x+4 }\] \[\Large \frac{ 3x+4 }{ (x-2)(3x+4) } - \frac{ 1 }{ 3x+4 }\] \[\Large \frac{ 3x+4 }{ (x-2)(3x+4) } - \frac{ 1(x-2) }{ (x-2)(3x+4) }\] \[\Large \frac{ 3x+4 }{ (x-2)(3x+4) } - \frac{ x-2 }{ (x-2)(3x+4) }\] \[\Large \frac{ 3x+4 -( x-2) }{ (x-2)(3x+4) }\] \[\Large \frac{ 3x+4 -x+2 }{ (x-2)(3x+4) }\] \[\Large \frac{ 2x+6 }{ (x-2)(3x+4) }\] So \[\Large \frac{ 1 }{ x-2 } - \frac{ 1 }{ 3x+4 }\] simplifies to \[\Large \frac{ 2x+6 }{ (x-2)(3x+4) }\]

OpenStudy (anonymous):

thats what I get

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

the questions says state the restrictions and simplify and the answer is the following and the restrictions I understand are -4/3, 2 whic I can see how they get

OpenStudy (anonymous):

I just figured it out

OpenStudy (anonymous):

I can take 2 out of each top item

jimthompson5910 (jim_thompson5910):

yes you can factor it out if you want

jimthompson5910 (jim_thompson5910):

and the restrictions are in place to avoid division by zero

OpenStudy (anonymous):

thank you it took working with someone to me to see it. I can't believe i didn't see it earlier. thanks

OpenStudy (anonymous):

@jim_thompson5910 I can always rely on you thanks

jimthompson5910 (jim_thompson5910):

you're welcome, glad to be of help